OpenStudy (anonymous):

Determine the nature of the root : (x – 2a)(x – 2b) = 4ab

OpenStudy (anonymous):

when i solved , i get, x^2-2x(a+b)=0 D = b^2-4ac= 4(a+b)>0 Hence, roots are real.. i just want to confirm my answer

OpenStudy (anonymous):

D = b^2-4ac= 4(a+b)^2>0

OpenStudy (raden):

agree :)

OpenStudy (anonymous):

thank u

OpenStudy (raden):

yw